Article View: rec.puzzles
Article #2534Re: Help me with MH puzzle, also, a new puzzle.
From: mikem
Date: Wed, 06 Jan 1993 21:33
Date: Wed, 06 Jan 1993 21:33
60 lines
3566 bytes
3566 bytes
In article <1i7cb6INNe0@mirror.digex.com> adric@access.digex.com (William Johnson) writes: wj> When playing Russian Roulette, each member has a revolver with 6 chambers wj> and 1 bullet (may have been 2 bullets, I don't think it changes the answer, wj> just the likelihood of death). Each member then fires the gun twice into wj> his/her own head. Members may spin the revolver before each shot. Which wj> option gives the best chance of survival: wj> A) Spin shoot spin shoot wj> B) Spin shoot shoot wj> C) Shoot spin shoot wj> D) Shoot shoot. wj> In article <1i8at3INNjbl@iskut.ucs.ubc.ca> leon@unixg.ubc.ca (Leon ter Beek) writes: ltb> A=C if you didn't spin before you start the next round. Wait a minute. What's this about starting another round? The problem states, "Each member then fires the gun twice into his/her own head." Twice. Two shots. That's it. Nothing more. Furthermore, each member has his/her *own* *gun*, so the result of any other player's move is irrelevant too, since you use your own gun, not the gun just fired by another player. ltb> Spinning makes the events independent, ie. next shot same probability as ltb> first shot if you survive the first one. I'll go along with that. ltb> A: P(survive) = 5/6*5/6= 25/36 ltb> B: P(survive) = 5/6*4/6= 20/36 Second shot only 4 empty chamber left. ltb> C: P(survive) = 3/6*3/6= 9/36 Don't spin after preceding round. ltb> D: P(survive) = 3/6*2/6= 6/36 Whoa! Hold on! Why is first probability in C and D only 3/6? What assumption are you making about the state of the gun prior to the first shot? What is the basis for your making this assumption? wj> Ok, it seems obvious to me that the two options where you shoot twice wj> without a spin in the middle are worst. Because assuming you survive wj> the first shot (because if you didn't, it's over), you have now reduced wj> your chances of surviving the next chamber. Makes perfect sense to me. wj> But I don't see the difference between A and C, unless in case C you know wj> in advance that where you put the bullet either was or was not in the wj> next chamber to be fired. This may have been the answer, but I don't think wj> so, because reading their answer sheet, the probabilities weren't as clear wj> cut as that. Any suggestions? The missing piece of information in this puzzle is whether you know anything about the state of the gun before you choose an option. But the puzzle does not fill in this missing information. It's therefore not complete, and you can't answer it based on the information provided. Bad puzzle! Bad, bad puzzle! No biscuit! ------------------------------------------------------------------------------- Michael Maciolek - mikem@ll.mit.edu --- VOICE (617)981-3174 - FAX (617)981-0189 Network Engineer - MIT Lincoln Laboratory - Computer Telecommunications Systems ------------------------------------------------------------------------------- "It is not an optical illusion, it just looks like one."
Technical Details
Message-ID:
<1993Jan6.213351.29174@ll.mit.edu>Path:
arc1-ca.pugleaf.net!sparky!uunet!spool.mu.edu!agate!stanford.edu!rutgers!micro-heart-of-gold.mit.edu!xn.ll.mit.edu!ll.mit.edu!newsReferences:
<1992Dec18.192640.28521@CSD-NewsHost.Stanford.EDU> <1i7cb6INNe0@mirror.digex.com> <1i8at3INNjbl@iskut.ucs.ubc.ca>Show Full Headers
Complete Article Headers
Path: arc1-ca.pugleaf.net!sparky!uunet!spool.mu.edu!agate!stanford.edu!rutgers!micro-heart-of-gold.mit.edu!xn.ll.mit.edu!ll.mit.edu!news From: mikem@ll.mit.edu Newsgroups: rec.puzzles Subject: Re: Help me with MH puzzle, also, a new puzzle. Message-ID: <1993Jan6.213351.29174@ll.mit.edu> Date: 6 Jan 93 21:33:51 GMT References: <1992Dec18.192640.28521@CSD-NewsHost.Stanford.EDU> <1i7cb6INNe0@mirror.digex.com> <1i8at3INNjbl@iskut.ucs.ubc.ca> Sender: mikem@ll.mit.edu Organization: MIT Lincoln Laboratory Lines: 60